### v1

# ACING: Actor-Critic for Instruction Learning in Black-Box Large Language Models

Abstract
The effectiveness of Large Language Models (LLMs) in solving tasks vastly depends on the quality of the instructions, which often require fine-tuning through extensive human effort. This highlights the need for automated instruction optimization; however, this optimization is particularly challenging when dealing with black-box LLMs, where model parameters and gradients remain inaccessible.
We propose ACING, a task-specific prompt optimization approach framed as a stateless continuous-action Reinforcement Learning (RL) problem, known as the continuum bandit setting. ACING leverages an actor-critic-based method to optimize prompts, learning from non-differentiable reward signals. We validate ACING by optimizing prompts for ChatGPT on 30 instruction-based tasks, as well as a summarization task. ACING consistently outperforms baseline methods, achieving a median score improvement of 10 percentage points compared to the best baseline considered. Furthermore, ACING not only recovers but also surpasses human-crafted expert instructions, achieving up to a 39 percentage point improvement over human benchmarks.
ACING: Actor-Critic for Instruction Learning
in Black-Box Large Language Models
[…]
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable performance across a wide range of tasks Zhao et al. (2024); Touvron et al. (2023). This success is largely attributed to their strong instruction-following capabilities, which enable adaptation to diverse downstream applications Chen et al. (2023); Liu et al. (2023). These instructions, commonly referred to as prompts, play a crucial role in the performance of LLMs Wei et al. (2022); Zhu et al. (2024); Liu et al. (2023). Given the importance of prompts, researchers are increasingly interested in automating their optimization to reduce the need for manual adjustments, which is often a labor-intensive and costly process Reynolds and McDonell (2021); Mishra et al. (2021). As a result, developing efficient methods for automatic prompt optimization has become highly important to maximize LLMs performance.
Several methods for automated prompt optimization have been proposed in the literature. Some focus on optimizing continuous prompts, known as soft prompts Zhong et al. (2021); Li and Liang (2021), which are typically fed into a LLM after its embedding layer since they do not represent meaningful language that the LLM can interpret directly. Other methods aim at optimizing discrete prompts, known as hard prompts, by either generating diverse prompts Zhou et al. (2023) or refining existing ones Pryzant et al. (2023). For both soft and hard prompts, some approaches rely on gradient-based methods Shin et al. (2020); Lester et al. (2021); Li and Liang (2021). However, these methods require access to the gradients of the LLMs and are therefore limited to white-box LLMs. Meanwhile, the most powerful LLMs today are typically black-box models (e.g., ChatGPT OpenAI (2023a) and GPT-4 OpenAI (2023b)).
To overcome these challenges, recent studies have introduced gradient-free techniques for optimizing prompts in black-box LLMs Zhou et al. (2023); Prasad et al. (2023); Pryzant et al. (2023). While promising, these approaches rely on heuristic local search methods and fail to fully utilize the observation history, including previously tested instructions and their outcomes, when selecting new instructions to test. Consequently, they are less effective at balancing the inherent exploration-exploitation trade-off, leading to inefficient querying. This inefficiency becomes problematic and impractical when API calls to black-box LLMs involve significant time and financial costs.
A recent line of work has focused on optimizing instructions for black-box LLMs by searching through soft prompts and transforming them into discrete prompts using white-box LLMs. InstructZero Chen et al. (2024) optimizes discrete prompts for ChatGPT by applying Bayesian optimization over soft prompts. Building on the same line, INSTINCT Lin et al. (2024) replaces Bayesian optimization with NeuralUCB Zhou et al. (2020), a critic algorithm, to select actions, from a finite pool of soft prompts, initially generated with Sobol Eriksson et al. (2019).
Inspired by the success of actor-critic methods for control problems Haarnoja et al. (2018, 2019) and generative adversarial networks Goodfellow et al. (2014), our approach combines the hybrid white-box/black-box framework with an actor-critic Reinforcement Learning (RL) algorithm, operating in a stateless, infinite continuous-action space, and constrained by a strict budget of API calls. The actor aims to maximize the critic’s value while remaining as stochastic as possible to enhance exploration and simultaneously exploit promising actions. This approach increases the likelihood of finding optimal prompts and enables more efficient use of the limited number of API calls. ACING consistently outperforms baseline methods, achieving a median score improvement of 10 percentage points. Furthermore, our optimized prompts not only recover but also surpass human-crafted expert instructions, achieving up to a 39 percentage point improvement against human benchmarks.
An implementation of ACING is available at https://github.com/salmakh1/ACING.
2 Problem Formulation
2.1 Prompt Learning Objective
We aim to improve the performance of a black-box LLM, denoted by $f$, which can only be accessed through its API, while its internal parameters remain unknown. Given a task represented by an (unknown) distribution $(x,y)\sim\mathcal{D}$—where $x$ denotes possible inputs and $y$ the corresponding correct outputs—our goal is to find the optimal prompt $\tau^{\star}$ that maximizes the likelihood of $f$ producing correct outputs for a given task. This is evaluated using a scoring function $q(\cdot,\cdot)\in[0,1]$. The black-box model $f$ processes an input formed by concatenating ($\oplus$) the prompt $\tau$ with the sentence $x$, producing a predicted output $\hat{y}=f(\tau\oplus x)$. More formally, the objective is to maximize the expected score of the LLM in solving the task $\mathcal{D}$, defined as:
$\max_{\tau}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau\oplus x))\right].$ (1)
To achieve this, we utilize a validation dataset $\mathcal{V}=\{(x_{j},y_{j})\}_{j=1}^{m}$, where each pair consists of an input sentence $x_{j}$ and its corresponding ground truth output $y_{j}$. Our objective is to find the prompt that maximizes the scoring function $q(\cdot,\cdot)$ across the validation dataset, where $q(\hat{y}_{j},y_{j})$ measures the quality of the predicted output $\hat{y}_{j}$ against the true output $y_{j}$. Thus, our objective becomes finding the prompt $\tau^{\star}$ that maximizes the average score over the validation set $\mathcal{V}$. The derived prompt $\tau^{\star}$ is then evaluated on a separate test set $\mathcal{T}=\{(x^{\prime}_{j},y^{\prime}_{j})\}_{j=1}^{m^{\prime}}$ to assess its generalization performance.
2.2 Learning over a Continuous Space
Directly optimizing the prompt $\tau$ for the black-box LLM model $f$ presents substantial challenges due to the discrete combinatorial nature of token selection in $\tau$. To mitigate this challenge, prior approaches like InstructZero Chen et al. (2024) and INSTINCT Lin et al. (2024) employ a white-box LLM, represented by $h$, to convert the discrete optimization problem into a continuous one by introducing a soft prompt vector $\mathbf{z}\in\mathbb{R}^{d}$, which is a continuous $d$-dimensional vector representing the token embedding of a set of virtual tokens.
The white-box model $h$ has access to a dataset of exemplars, =$\mathcal{E}=\{(u_{j},v_{j})\}_{j=1}^{k}$, where each pair $(u_{j},v_{j})$ defines input-output text sequences that exemplify a downstream task. The white-box LLM $h$ serves as a proxy, mapping the soft prompt $\mathbf{z}$ into a discrete prompt $\tau$ suitable for input to the black-box LLM. Specifically, given the exemplars $\mathcal{E}$ and the vector $\mathbf{z}$, their concatenation is input to the white-box LLM, generating the discrete prompt $\tau(\mathbf{z})=h(\mathbf{z},\mathcal{E})$. This generated prompt $\tau(\mathbf{z})$ is prepended to a test input $x_{j}$ from the validation set $\mathcal{V}$, and the combined input is provided to the black-box LLM $f$ to generate an output $\hat{y}_{j}=f(\tau(\mathbf{z})\oplus x_{j})$. The output $\hat{y}_{j}$ is then evaluated against the true output $y_{j}$ using the scoring function $q(\hat{y}_{j},y_{j})$. By using a fixed set of exemplars $\mathcal{E}$, the original discrete problem (Equation (1)) of finding the optimal prompt $\tau$ is effectively transformed into a continuous optimization problem over the soft prompt vector $\mathbf{z}$, as follows:
$\max_{\mathbf{z}\in\mathbb{R}^{d}}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau(\mathbf{z})\oplus x))\right].$ (2)
2.3 Reducing the Continuous Space
The soft prompt $\mathbf{z}$ is typically high-dimensional (e.g., $d=5120\times N_{z}$, with $N_{z}$ the number of virtual tokens, when $h$ is Vicuna-13B Chiang et al. (2023)), which makes it a challenging problem. Therefore, we employ random projection techniques to reduce the input dimension as done in prior works Chen et al. (2024); Lin et al. (2024). Specifically, given a matrix $P\in\mathbb{R}^{d\times d^{\prime}}$ where $d^{\prime}\ll d$, with randomly sampled elements, and a continuous vector $\mathbf{a}\in\mathbb{R}^{d^{\prime}}$, the vector $\mathbf{z}=P\mathbf{a}$ is used as the soft prompt. By substituting $\mathbf{z}$ with $P\mathbf{a}$, the input variable to be optimized in Eq. 2 becomes $\mathbf{a}$, thereby reducing the input dimension of the optimization problem to $d^{\prime}$:
$\max_{\mathbf{a}\in\mathbb{R}^{d^{{}^{\prime}}}}\quad\mathbb{E}_{(x,y)\sim\mathcal{D}}\left[q(y,f(\tau(P\mathbf{a})\oplus x))\right].$ (3)
For fairness with previous works, the reduced input dimension $d^{\prime}$, referred to as the intrinsic dimension, is chosen as $d^{\prime}=10$ Chen et al. (2024); Lin et al. (2024). An ablation study on the intrinsic dimension is discussed in § 5.4.3. Furthermore, to maintain consistency with these prior studies, we similarly use a projection matrix $P$ with elements uniformly and independently sampled from $\text{Uniform}(-1,1)$, and bound the vector such that $\mathbf{a}\in[0,1]^{d^{\prime}}$.
3 Actor-Critic Approach for Instruction Learning in Black-Box LLMs
3.1 Instruction Learning as an RL Problem
In this work, we frame the prompt-learning problem as an agent navigating an infinite action space within a $d^{{}^{\prime}}$-dimensional unit hypercube, searching within an infinite number of actions $\mathbf{a}\in[0,1]^{d^{{}^{\prime}}}$, in contrast to previous works that consider a finite number of actions initially generated with Sobol Eriksson et al. (2019); Lin et al. (2024). As part of the environment, these continuous actions are projected into a higher-dimensional space (§ 2.3) and then transformed into discrete prompts, through a white-box LLM (§ 4). These discrete prompts are evaluated by a black-box LLM over a validation dataset, after which a reward is computed and returned to the agent. We consider a finite horizon $T$, representing the maximum number of actions played (i.e., the budget of evaluated instructions), after which the agent ceases exploration. The agent’s goal is to discover the optimal continuous action, $\mathbf{a}^{\star}$, which produces the optimal discrete prompt and the highest reward.
Unlike in general RL settings, where the agent encounters multiple states with varying reward distributions, the agent we consider operates in a stateless, stochastic environment, commonly referred to as a stochastic bandit setting Slivkins (2019); Lattimore and Szepesvári (2020). In this context, at each iteration $t\leq T$ (where $t$ is incremented with every black-box LLM call), the agent’s action $\mathbf{a}_{t}$ generate stochastic rewards, $r_{t}\in[0,1]$, without affecting the reward distributions associated with those actions. The prompt-learning objective is modeled as a stochastic best-arm identification problem in a stateless, continuous action space RL, specifically in an infinite continuum bandit setting.
Algorithm 1 Stateless Soft Actor-Critic
0:  Evaluation budget $T$
1:  Initialize $\boldsymbol{\theta}$, $\mathbf{w}$ (e.g., randomly)
2:  for $t\leftarrow 1$ to $T$ do
3:     Choose action $\mathbf{a}_{t}\sim\pi(\cdot\mid\boldsymbol{\theta})$
4:     Take action $\mathbf{a}_{t}$ and observe reward $r_{t}$
5:     $\mathbf{w}\leftarrow\mathbf{w}-\lambda\hat{\nabla}\mathbf{J}_{Q}(\mathbf{w})\quad$ (Equation (5))
6:     $\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}+\beta\hat{\nabla}\mathbf{J}_{\pi}(\boldsymbol{\theta})\quad$ (Equation (6))
7:     $\alpha\leftarrow\alpha-\gamma\hat{\nabla}\mathbf{J}_{\alpha}(\alpha)\quad$ (Equation (7))
8:  end for
3.2 Actor-Critic Continuum Bandits
RL techniques are generally used to learn optimal policies and actions leading to best rewards. These techniques can be classified into value-based (critic-centric), policy-gradient (actor-centric), or hybrid approaches. Actor-critic algorithms merge the benefits of both by employing two key components: the actor, responsible for policy optimization, and the critic, which evaluates the policy (Sutton and Barto, 2018). This hybrid approach has demonstrated success in various domains, including continuous control problems Mnih et al. (2016); Haarnoja et al. (2018, 2019) and with generative adversarial networks (GANs) (Goodfellow et al., 2014), where a generator and a critic network collaboratively optimize performance.
Figure 1: Pipeline of ACING. In each iteration, a soft prompt along with several examples of the target task are provided to the white-box LLM to generate an instruction. This instruction is then used to query the black-box LLM, which produces answers to the target task queries. The resulting score is returned to the agent as a reward, which is used to update its networks and adjust its policy. Both LLMs remain frozen throughout the process.
Inspired by the success of actor-critic approaches, we introduce a (stateless) actor-critic algorithm, as provided in Algorithm 1, tailored to our infinite continuum bandit setting, enabling autonomous learning of effective prompts with a constrained evaluation budget and outperforming previous state-of-the-art black-box prompt learning approaches. Both the actor and the critic are parameterized by neural networks. In typical RL problems, the agent interacts with an environment with varying states, so its policy depends on the current state. That is, the policy network generates actions based on the current state $s$, $\pi(.|s;\theta)$, where $\theta$ denotes the parameters. However, in the problem we consider, the setting is stateless. Therefore, we define a policy network $\pi(.;\theta)$ with a constant input that produces a $d^{{}^{\prime}}$-dimensional vector as output $\mathbf{a}$, representing the continuous action. For the critic, we use a neural network with a $d^{{}^{\prime}}$-dimensional input (the action) and a single output that estimates the action’s quality, thereby assessing the quality of the policy that proposed it. The critic network’s parameters are trained to minimize the following:
$\min_{\mathbf{w}}J_{Q}(\mathbf{w})\triangleq\mathbb{E}_{\mathbf{a}\sim\mathcal{D}}\left[\frac{1}{2}\left(Q_{\mathbf{w}}\left(\mathbf{a}\right)-r\left(\mathbf{a}\right)\right)^{2}\right],$ (4)
which can be optimized with stochastic gradients
^$\hat{\nabla}_{\theta}J_{Q}(\mathbf{w})=\nabla_{\mathbf{w}}Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)\left(Q_{\mathbf{w}}\left(\mathbf{a}_{t}\right)-r\left(\mathbf{a}_{t}\right)\right),$ (5)
where $Q$ is the critic-network with parameters $\mathbf{w}$. The actor-critic learning process alternates between policy-network and critic-network updates.
3.3 Enhancing Exploration with Entropy
In general, RL learning approaches aim to maximize the expected cumulative rewards up to a horizon $T$, expressed as $\max_{\pi(\cdot)}\sum_{t=1}^{T}\mathbb{E}_{a_{t}\sim\pi(\cdot)}\left[r(a_{t})\right].$ In contrast, black-box prompt learning, framed as a best-arm identification problem, seeks the best prompt without necessarily maximizing the expected cumulative rewards. To maximize the cumulative rewards, the agent must manage the exploration-exploitation trade-off. Conversely, for the best-arm identification problem, the agent should explore as extensively as possible within the $T$ evaluation budget to potentially identify the best action, leading to a pure exploration setting. Although it might seem that the agent should act purely at random to find the best action, some degree of exploitation may still be beneficial to guide the agent toward the optimal choice.
To achieve a high exploration rate while exploiting prior data, we consider the maximum entropy objective, as discussed in previous RL works Ziebart (2010); Haarnoja et al. (2018). This objective promotes stochastic policies by augmenting the reward with an entropy term, expressed as:
$\min_{\theta}J_{\pi}(\theta)\triangleq\mathbb{E}_{\mathbf{a}\sim\pi(.;\theta)}\left[\alpha\log\left(\pi\left(\mathbf{a};\theta\right)\right)-Q_{\mathbf{w}}\left(\mathbf{a})\right)\right],$ (6)
where the temperature $\alpha$ controls the stochasticity of the agent. The above objective is a simplification of the objective in the Soft Actor-Critic (SAC) algorithm Haarnoja et al. (2018), eliminating the state, as we consider a stateless environment and focus on instantaneous rather than cumulative rewards. The term $-\mathbb{E}_{a\sim\pi(.;\theta)}\left[\log\pi(a)\right]$ represents the entropy of the policy, which we seek to maximize alongside the critic values, as entropy maximization is crucial for encouraging the actor to explore different actions. A high $\alpha$ encourages exploration, whereas a lower value of $\alpha$ favors exploitation.
Furthermore, to avoid fine-tuning, and to achieve adaptive exploration, the entropy coefficient $\alpha$ can be adjusted dynamically to maintain a target entropy level Haarnoja et al. (2019), $H_{\text{target}}$, ensuring sufficient exploration, as follows:
$\min_{\alpha}\mathbf{J}_{\alpha}\triangleq-\mathbb{E}_{a\sim\pi(.)}\left[\alpha\cdot\left(\log\pi(\mathbf{a};\theta)+H_{\text{target}}\right)\right].$ (7)
We use a stochastic approximation of gradients, specifically the Adam algorithm Kingma and Ba (2017); Reddi et al. (2018), to minimize the losses via stochastic gradient descent. The learning rates $\lambda$, $\beta$, and $\gamma$ are used for the Q-network, the policy networks, and the temperature, respectively.
Figure 2: Illustration of the prompt generation and testing inside the environment.
4 Methodology
The methodology of our proposed algorithm, ACING, is illustrated in Fig. 1, showing the actor-critic interaction with the environment. In Fig. 2, the environment is zoomed in, using the larger_animal dataset as an example. In the following, we provide a detailed explanation of the full methodology.
Overview. In each iteration $t\leq T$, the actor-critic agent generates a continuous vector “action” $\mathbf{a}$ (step 1). This action is then played on the environment, which projects $\mathbf{a}$ into the appropriate space using a fixed projection matrix $P$ to obtain $\mathbf{z}$. The environment then concatenates the projected vector $\mathbf{z}$ with a set of exemplars’ embeddings from $\mathcal{E}$ and feeds it into a white-box LLM $h$ (step 2). The white-box LLM produces a discrete prompt, $\tau$, which is evaluated using the validation dataset $\mathcal{V}$ based on the responses from the black-box LLM $f$ (step 3). The black-box LLM generates a prediction, which is then contrasted to the true labels of the validation examples, and a score function provides a final reward to the critic. This reward is used to compute the critic’s loss and update both the critic and actor networks accordingly.
[…]
By repeating steps 1 to 4 until the allowed budget $T$ is exhausted, the agent returns the best instruction prompt, denoted as $\tau^{\star}$. This prompt is then evaluated using the test set $\mathcal{T}$ on the black-box LLM, yielding the final score $\mathcal{S}^{\star}=\frac{1}{m^{\prime}}\sum_{i=1}^{m^{\prime}}q(f(\tau^{\star},x^{\prime}_{i}),y^{\prime}_{i})$ (using the evaluation template depicted in Fig. 1 bottom right).
Category Task Human Instruction ACING Instruction (Ours) Human Score Our Score
Spelling Second_word_letter Extract the second letter of the input word. Output the first letter of each word that is a vowel (a, e, i, o, u) Example: Input: year Output: e Input: trust Output: r Input: qualification Output: u Input: defendant Output: e 0.96 0.59
Syntax Negation Negate the input sentence. Flip the truth value of the statements in the input 0.81 0.82
Lexical Semantics Antonyms Write a word that means the opposite of the input word. Take a word and change it to its opposite 0.70 0.82
Synonyms Write a word with a similar meaning to the input word. Input a word that is a synonym for the word that was output 0.14 0.13
Phonetics Rhymes Write a word that rhymes with the input word. Input the word that the program thought I was inputting and then output the word that program thought I was inputting 0.61 1.00
Semantics Cause_and_effect Find which of the two given cause and effect sentences is the cause. Find the sentence that is the cause of the effect in the pair of sentences 0.97 0.92
Style Informal_to_formal Rephrase the sentence in formal language. Convert the input into output using the same word order and with the same meaning 0.63 0.5
Multi- lingual Translation_en-de Translate the word into German. Provide a translation for each word in the English text into German 0.81 0.84
Translation_en-es Translate the word into Spanish. Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 0.89 0.88
Translation_en-fr Translate the word into French. Create a program that would take an English word as input and output its French equivalent 0.86 0.87
GLUE Sentiment Determine whether a movie review is positive or negative. Classify each input as positive or negative based on the assessment of the corresponding movie 0.89 0.90
Sentence_similarity Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly. Find a sentence pair that is probably not similar, and the output is 3 - probably 0.00 0.21
median score 0.81 0.83
average score 0.69 0.71
# best-performing tasks 5 7
Table 1: Tasks from the instruction-induction datasets where the human and ACING test scores differed. For each task, we provide the corresponding human instruction as proposed in Honovich et al. (2023) and our best discovered instruction. We tested these instructions on the test dataset and report their respective scores.
Category Task APE InstructZero INSTINCT ACING
Spelling Letters_list 0.59 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00)
First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.00 (0.00) 0.35 (0.09) 0.39 (0.28) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00)
Negation 0.79 (0.00) 0.65 (0.10) 0.58 (0.22) 0.71 (0.06)
Lexical Semantics Antonyms 0.79 (0.02) 0.76 (0.00) 0.84 (0.01) 0.74 (0.01)
Synonyms 0.14 (0.01) 0.22 (0.11) 0.19 (0.08) 0.13 (0.02)
Word_unscrambling 0.54 (0.00) 0.59 (0.06) 0.54 (0.02) 0.50 (0.07)
Phonetics Rhymes 0.59 (0.01) 0.99 (0.01) 0.36 (0.04) 0.57 (0.31)
Numerical Sum 0.87 (0.01) 1.00 (0.00) 0.70 (0.21) 0.69 (0.31)
Diff 0.00 (0.00) 1.00 (0.00) 0.93 (0.09) 1.00 (0.00)
Knowledge Larger_animal 0.72 (0.02) 0.63 (0.07) 0.81 (0.09) 0.84 (0.07)
Periodic_elements 0.99 (0.01) 0.96 (0.03) 1.00 (0.00) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.52 (0.09) 0.55 (0.11) 0.69 (0.15)
Common_concept 0.03 (0.02) 0.14 (0.04) 0.09 (0.04) 0.19 (0.05)
Object_counting 0.30 (0.02) 0.38 (0.06) 0.40 (0.12) 0.41 (0.03)
Odd_one_out 0.32 (0.02) 0.57 (0.02) 0.25 (0.18) 0.64 (0.00)
Orthography_starts_with 0.23 (0.01) 0.41 (0.09) 0.54 (0.06) 0.60 (0.12)
Taxonomy_animal 0.02 (0.02) 0.67 (0.14) 0.85 (0.06) 0.71 (0.02)
Auto_categorization 0.31 (0.01) 0.29 (0.02) 0.07 (0.07) 0.29 (0.04)
Word_sorting 0.58 (0.01) 0.64 (0.05) 0.23 (0.20) 0.70 (0.03)
CLUE Sentence_similarity 0.00 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.07)
Sentiment 0.90 (0.00) 0.88 (0.03) 0.88 (0.02) 0.89 (0.01)
Translation Num_to_verbal 0.13 (0.02) 0.99 (0.01) 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.83 (0.01) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01)
Translation_en-es 0.86 (0.01) 0.67 (0.24) 0.89 (0.00) 0.87 (0.02)
Translation_en-fr 0.88 (0.01) 0.77 (0.06) 0.85 (0.02) 0.83 (0.01)
Style Informal_to_formal 0.57 (0.01) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.25 (0.00)
median score 0.49 0.66 0.64 0.71
# best-performing tasks 9 8 7 14
Table 2: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, and INSTINCT. The bottom rows report the median score and total number of best-performing tasks for each method.
5 Experiments
We tackle instruction learning for ChatGPT. We conduct instruction induction tasks using 30 datasets from Honovich et al. (2023); Chen et al. (2024), along with a data summarization task from Gliwa et al. (2019). Furthermore, we compare the best-learned instructions from our ACING algorithm against human-provided instructions from Honovich et al. (2023), as shown in Table 1, with additional best-learned instructions in Table 9. Moreover, we compare ACING with three recent representative baselines for black-box instruction learning: APE (Zhou et al., 2023), InstructZero (Chen et al., 2024), and INSTINCT Lin et al. (2024), with results provided in Table 2.
Following previous works and to ensure fairness, we do not fine-tune the budget. Instead, like previous works, we consider a fixed budget of black-box LLM evaluations, set to $T=165$ Chen et al. (2024); Lin et al. (2024). Moreover, like previous works, we use Vicuna-13B Chiang et al. (2023) as the white-box LLM $h$. Furthermore, for previous works, we use their default (tuned) hyperparameters, including the intrinsic dimension $d^{\prime}=10$ and the number of soft tokens $N_{z}=5$. For fairness, we refrain from fine-tuning these parameters for our method and use the same values as in prior works. This ensures that our ACING algorithm searches in the same space, $[0,1]^{10}$, and uses the same total number of queries to the black-box LLM as APE, InstructZero, and INSTINCT for a fair comparison. For each algorithm, after identifying the best instruction using the validation set $\mathcal{V}$, we evaluate the discovered instruction on a separate test set $\mathcal{T}$ and report the test accuracy as the score.
We further conduct ablation studies on the key design choices for ACING, including the use of small learning budgets, providing reward plots for various selected tasks. Furthermore, we test the impact of using different white-box language models, intrinsic dimensions, and numbers of exemplars. More details on the experiments are provided in the Appendix.
5.1 ACING vs. Humans
We use human instructions provided in Honovich et al. (2023) for various instruction-induction tasks, test them, and compare their scores against those of our best instructions. The complete list of selected tasks is provided in Table 1 in the Appendix. Several instructions were solved by both human experts and our approach, i.e., achieving a perfect test score of 1.00. For clarity and brevity, we omit these results from the main paper and focus only on tasks where the scores of humans and our method differed, as shown in Table 1.
As shown in these tables, our approach outperforms human instructions in several tasks, sometimes by a significant margin. For example, in the rhyming task, our method achieved a 0.39 improvement in the test score. The human instruction for this task—Write a word that rhymes with the input word.—only resulted in a score of 0.61. In contrast, the best ACING instruction, which was an intriguing prompt: Input the word that the program thought I was inputting and then output the word that program thought I was inputting, achieved a perfect score of 1.00.
Furthermore, in the challenging sentence similarity task, while the human-provided instruction achieved a score of 0.00, our approach improved the test score by up to 0.21. However, the instruction yielding the highest score appeared to bias the prediction towards outputting “3” due to the following instruction: Find a sentence pair that is probably not similar, and the output is 3 - probably, which resulted in the best score. The second-best instruction, proposed by ACING, achieved an improvement of 0.14 over the human score and was as follows: Generate a probability score for each pair of sentences, where 1 - probably not means that the two sentences are not related, 0 - definitely not means that the two sentences are not related, and 3 - probably means that the two sentences are related. Overall, our approach recovers several human test scores while being fully automatic, and in many cases surpasses human performance in terms of the number of successful tasks, average score, and median score across tasks. For the curious reader, we provide all our best instructions along with their test scores in Table 10 in the Appendix.
5.2 ACING vs. Other Optimization Methods
We compare our method against recent baselines on the 30 instruction-induction datasets. Table 2 presents the average test accuracy (along with the standard deviation) over three independent runs, using three different seeds. For each seed, we selected the best instruction achieved by each method and evaluated it on the testing dataset. The table demonstrates that our method, ACING, outperforms others by achieving the highest accuracy in 14 out of 30 tasks, compared to INSTINCT, InstructZERO, and APE which succeeded in 9 tasks or less each. Additionally, ACING achieves the highest median accuracy across tasks, with a value of 0.71, which is  22 percentage points higher than APE. Further details, including the best prompt achieved for each task and corresponding test scores, are in the Appendix, Table 10. Moreover, we compare the performance of ACING against other methods on a summarization task using SAMSum dataset Gliwa et al. (2019) in Table 3, which shows that ACING outperforms the other methods.
In the following section, we demonstrate that the median score of ACING can be further improved by 5 percentage points using a simple two-phase exploration technique (splitting the budget), resulting in a 10-percentage-point improvement over the best baseline considered. Furthermore, in ablation studies, we show that fine-tuning the intrinsic dimensions $d^{{}^{\prime}}$ can lead to additional improvements.
Metric APE InstructZero INSTINCT ACING
ROUGE-1 0.35 (0.01) 0.33 (0.00) 0.36 (0.01) 0.37 (0.01)
ROUGE-2 0.12 (0.00) 0.11 (0.00) 0.14 (0.00) 0.14 (0.00)
ROUGE-L 0.25 (0.00) 0.24 (0.01) 0.27 (0.01) 0.28 (0.01)
Table 3: Average test performance (and standard deviations) for instruction optimization for summarization task using SAMSum dataset.
5.3 Splitting the Budget
Due to the stochastic nature of the black-box LLM, the same instruction may yield different rewards when evaluated by the LLM. For more robust prompt selection, the budget $T$ can be split into two parts: an actor-critic exploration phase (repeating steps 1 to 4), followed by a uniform exploration phase. In the latter phase, the top $p$ prompts discovered by the actor-critic are evaluated multiple times, $k$ times each, using the black-box LLM. Specifically, the first phase uses $T-p\cdot k$ API calls, while the second uses the remaining calls. The prompt with the highest average reward across repetitions can then be used at test time. In Table 5 in the Appendix, we show that ACING, using this budget split technique with an initial exploration budget of $150$ and a remaining budget of $15$ calls (allocated to uniformly explore the top $p=5$ prompts, each evaluated $k=3$ times), achieves improved median scores across tasks, resulting in 10 percentage points improvement over the best baseline considered and higher test accuracy on 15 tasks compared to other methods.
5.4 Ablation Studies
5.4.1 Using Smaller Budgets
In the main paper, we report the final test score after a fixed budget of 165 black-box LLM calls, mainly following previous work Lin et al. (2024); Chen et al. (2024), avoiding any potential advantage that could arise from optimizing this number. In the Appendix, we provide reward plots for the ACING approach, showing the best achieved reward over the calls. As shown in various plots in Figure 3, the ACING approach found the optimal instruction (achieving a reward of 1) within just a few black-box calls. Some tasks required fewer than 10 API calls to find the optimal instruction, as in the case of ‘active to passive’ and ‘letters list’, and fewer than 20 calls for tasks like ‘translation’ and ‘diff’. It can be seen that the vast majority of tasks achieved their best reward value within the first 60 to 80 calls, demonstrating that ACING can even be used for more constrained budgets.
5.4.2 Using a Different White-box Model
In the main paper, we present results using Vicuna as the white-box language model, following the setup of prior work. To evaluate the impact of different white-box models on the performance of our method, we also tested ACING with WizardLM-13B-v1.2 Xu et al. (2024) compared to Vicuna-13B-v1.3. The experiments in Table 8 indicate that WizardLM achieved a higher median test score across all tasks and excelled in a greater number of top-performing tasks, demonstrating that tuning the choice of the white-box model can improve results.
5.4.3 Using Different Action Dimensionalities
In the main paper, we present results using actions with dimension $d^{\prime}=10$, following the setup of prior work. To evaluate the performance of ACING with different dimensionalities, we conducted an experiment using $d^{\prime}\in\{5,10,20,40,100\}$, keeping other parameters fixed, for a budget of $165$ calls. The results, shown in Table 6, indicate that while the smallest dimension, $d^{\prime}=5$, recovered the best scores for some tasks, it generally has the lowest performance across most tasks. Furthermore, both $d^{\prime}=10$ and $d^{\prime}=20$ yield similar performance in terms of the number of best-performing tasks, indicating a low sensitivity to this parameter. For the larger dimension, $d^{\prime}=40$, the method achieved the highest number of best-performing tasks, showing that tuning (increasing) $d^{\prime}$ can improve results.
5.4.4 Using a Different Number of Exemplars
In the main paper, we present results using five exemplars, following the setup of prior work. To evaluate the performance of ACING with a smaller number of exemplars, we conducted an experiment using a single exemplar, i.e., $|\mathcal{E}|=1$. Results shown in Table 7 confirm that using five exemplars generally yield higher test scores and a greater number of best-performing tasks than a single exemplar. However, a single exemplar still performs well, matching the performance of five exemplars across several tasks (e.g., phonetics, summation, morpho-syntax, translation) and sometimes even outperforming them. Nonetheless, five exemplars offer an advantage, particularly for more cognitively complex tasks.
6 Related Work
Many early approaches such as AutoPrompt Shin et al. (2020), FluentPrompt Shi et al. (2023), and other soft prompt-based methods Lester et al. (2021); Li and Liang (2021); Zhong et al. (2021) focused on optimizing prompts for white-box LLMs and cannot be directly applied to black-box LLMs. Moreover, several techniques have also been proposed for optimizing prompts for grey-box LLMs: BBT Sun et al. (2022b), BBTv2 Sun et al. (2022a), and clip-tuning Chai et al. (2022) use evolutionary algorithms for this purpose. However, these methods still require access to input token embeddings and output logits, making them unsuitable for settings where only query access is allowed.
Other approaches Li and Liang (2021); Zhong et al. (2021) employ continuous soft prompts by fine-tuning parameters of specific input tokens, but these often generate human-unreadable prompts and are impractical for API-accessible LLMs. Some other methods Deng et al. (2022); Zhang et al. (2023), created for optimizing discrete prompts instead of continuous ones, can optimize discrete tokens but require access to the LLM’s output distribution, which is typically unavailable in black-box LLMs. Zhou et al. (2023) proposed generating candidate prompts through resampling without specific optimizations, while Guo et al. (2024) extended this model-free approach by using evolutionary algorithms. However, these methods usually require a large number of queries.
Instruction optimization in black-box LLMs recently focused on searching through the space of soft prompts, then relying on open-source white-box LLMs as a means to transform these soft prompts into discrete prompts. InstructZero Chen et al. (2024) learns discrete prompts for ChatGPT by searching through the continuous space of soft prompts using Bayesian optimization, which are then transformed into discrete prompts by a white-box LLM. INSTINCT Lin et al. (2024), a successor to this approach, relies on the same approach. However, instead of Bayesian optimization, it uses NeuralUCB Zhou et al. (2020), a contextual bandit (selection) algorithm that searches within a finite set of continuous actions generated using Sobol Eriksson et al. (2019). Our approach combines the hybrid strategy of using white-box and black-box LLMs with an actor-critic approach that operates in a stateless, infinite continuous-action space.
TRIPLE Shi et al. (2024) modeled the prompt learning problem as a best-arm-identification problem. However, it only considered the problem as a selection task, assuming a given pool of discrete prompts. In contrast, our method addresses both the generation and selection of prompts.
7 Conclusion
ACING presents a promising approach for learning prompts for black-box LLMs by framing it as a stateless continuous-action RL problem. Our method shows significant improvements over recent baseline methods. Evaluations across 31 datasets reveal that ACING enhances prompt quality, achieving a median score improvement of 10 percentage points and outperforming human-crafted instructions by up to 39 percentage points. This sets a new benchmark for automated instruction optimization in black-box LLMs.
8 Limitations
While our proposed results were achieved using default and commonly used values for hyperparameters in network design, learning rates, and following previous works for some common parameters, the results presented depend on these choices. These results could be further improved with fine-tuning, as demonstrated in the ablation studies (§ 5.4). Another limitation of this work, similar to previous studies, is the risk of overfitting when dealing with a limited number of validation examples. This can result in a prompt that minimizes the loss on the validation set but does not generalize well to the test dataset, leading to a discrepancy between training and test performance. Another limitation of this work, similar to previous work (Lin et al., 2024; Chen et al., 2024), is the reliance on a white-box model, where the choice of model may lead to performance variations, as demonstrated in § 5.4.2. Avoiding this limitation requires acting directly on the black-box LLM, which involves handling actions in a large discrete action space. This necessitates adapting current RL approaches for large discrete action spaces, as studied in Fourati et al. (2024) and the references therein.
Appendix A Experimental Details
A.1 Hyperparameters Remark
Across the diverse tasks, in the main paper, the same hyperparameters were used, which shows that the algorithm generalizes well across the 30 tasks without specifically tuning hyperparameters in each task. A summary of the key parameters cab be found in the following Table. An implementation of ACING is available at https://github.com/salmakh1/ACING.
Hyperparameter Choice
White-box $h$ Vicuna-13B
Actor-network $(1-1024-256-10)$
Critic-network $(10-128-128-1)$
Budget $T$ 165
Intrinsic (action) dimension $d^{\prime}$ 5
Number of soft tokens $N_{z}$ 5
Soft prompt dimension $d$ 5120 * $N_{z}$
Number of exemplars $|\mathcal{E}|$ 5
Number of tokens generated by Wb 64
Table 4: Key hyperparameters and their values.
A.2 Actor-Critic Details
Across all the tasks, we used three fully-connected layers for both the actor $(1-1024-256-10)$ and the critics $(10-128-128-1)$ networks, with learning rates fixed at $3\cdot 10^{-4}$ for each. We used two critic networks and consider their minimum as the actual critic. We learn the entropy parameter $\alpha$ using a learning rate of $9\cdot 10^{-4}$.
Appendix B ACING Rewards over the (Calls) Steps
In the main paper, we report the final test score after a fixed budget of 165 black-box LLM calls. In this section, we provide reward plots for the ACING approach, showing the best-achieved reward within the conducted calls. As shown in various plots in Figure 3, the ACING approach found the optimal prompt (achieving a reward of 1) within just a few black-box calls. Some tasks required fewer than 10 API calls to find the optimal instruction, such as for ‘active to passive’ and ‘letters list’, and fewer than 20 for tasks like ‘translation’ and ‘diff’. It can be seen that the vast majority of tasks achieved their best reward value within the first 60 to 80 calls, demonstrating that ACING can even be used for much more constrained budgets. The choice of 165 calls was mainly based on previous work Lin et al. (2024); Chen et al. (2024), avoiding any potential advantage that could come from optimizing this number.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Figure 3: Reward plots for running ACING on various selected tasks, showing the highest achieved reward on the y-axis until each API call (step), with the x-axis representing the number of API calls.
Appendix C ACING with Budget Splitting
Due to the stochastic nature of the black-box LLM, the same instruction may yield different rewards when evaluated by the LLM. To address this, we add a mechanism for more robust decision-making. The budget $T$ is split into two parts: an exploration phase where steps 1 to 4 are repeated, and an exploitation phase where the best $p$ prompts are evaluated multiple times, $k$ times each, using the black-box LLM. The exploration phase uses $T-p\cdot k$ API calls, with the remaining calls used for exploitation. Finally, the prompt with the highest average reward across repetitions is used at test time. In Table 5, we demonstrate that ACING, with an exploration budget of $T=150$ and the remaining 15 calls allocated to uniform exploration of the top $p=5$ prompts (evaluated $k=3$ times each), achieves improved median scores across tasks and higher test accuracy on 15 tasks compared to previous work, unlike without the splitting technique which achieved higher test accuracy on 14 tasks compared to the previous work.
Category Task APE InstructZero INSTINCT ACING150+15 ACING165
(budget splitting) (main paper)
Spelling Letters_list 0.59 (0.02) 1.00 (0.00) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00)
First_word_letter 0.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.00 (0.00) 0.35 (0.09) 0.39 (0.28) 0.40 (0.17) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.99 (0.01) 0.95 (0.03) 0.99 (0.01) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 0.98 (0.01) 1.00 (0.00) 1.00 (0.00) 1.00 (0.00)
Negation 0.79 (0.00) 0.65 (0.10) 0.58 (0.22) 0.82 (0.00) 0.71 (0.06)
Lexical Semantics Antonyms 0.79 (0.02) 0.76 (0.00) 0.84 (0.01) 0.76 (0.06) 0.74 (0.01)
Synonyms 0.14 (0.01) 0.22 (0.11) 0.19 (0.08) 0.12 (0.02) 0.13 (0.02)
Word_unscrambling 0.54 (0.00) 0.59 (0.06) 0.54 (0.02) 0.59 (0.05) 0.50 (0.07)
Phonetics Rhymes 0.59 (0.01) 0.99 (0.01) 0.36 (0.04) 0.60 (0.38) 0.57 (0.31)
Numerical Sum 0.87 (0.01) 1.00 (0.00) 0.70 (0.21) 0.81 (0.30) 0.69 (0.31)
Diff 0.00 (0.00) 1.00 (0.00) 0.93 (0.09) 0.97 (0.04) 1.00 (0.00)
Knowledge Larger_animal 0.72 (0.02) 0.63 (0.07) 0.81 (0.09) 0.86 (0.06) 0.84 (0.07)
Periodic_elements 0.99 (0.01) 0.96 (0.03) 1.00 (0.00) 0.98 (0.03) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.44 (0.09) 0.52 (0.09) 0.55 (0.11) 0.76 (0.18) 0.69 (0.15)
Common_concept 0.03 (0.02) 0.14 (0.04) 0.09 (0.04) 0.10 (0.01) 0.19 (0.05)
Object_counting 0.30 (0.02) 0.38 (0.06) 0.40 (0.12) 0.48 (0.11) 0.41 (0.03)
Odd_one_out 0.32 (0.02) 0.57 (0.02) 0.25 (0.18) 0.59 (0.05) 0.64 (0.00)
Orthography_starts_with 0.23 (0.01) 0.41 (0.09) 0.54 (0.06) 0.54 (0.15) 0.60 (0.12)
Taxonomy_animal 0.02 (0.02) 0.67 (0.14) 0.85 (0.06) 0.53 (0.34) 0.71 (0.02)
Auto_categorization 0.31 (0.01) 0.29 (0.02) 0.07 (0.07) 0.27 (0.06) 0.29 (0.04)
Word_sorting 0.58 (0.01) 0.64 (0.05) 0.23 (0.20) 0.72 (0.02) 0.70 (0.03)
CLUE Sentence_similarity 0.00 (0.00) 0.10 (0.00) 0.00 (0.00) 0.13 (0.08) 0.13 (0.07)
Sentiment 0.90 (0.00) 0.88 (0.03) 0.88 (0.02) 0.88 (0.03) 0.89 (0.01)
Translation Num_to_verbal 0.13 (0.02) 0.99 (0.01) 1.00 (0.00) 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.83 (0.01) 0.82 (0.01) 0.77 (0.02) 0.82 (0.01) 0.82 (0.01)
Translation_en-es 0.86 (0.01) 0.67 (0.24) 0.89 (0.00) 0.86 (0.02) 0.87 (0.02)
Translation_en-fr 0.88 (0.01) 0.77 (0.06) 0.85 (0.02) 0.85 (0.02) 0.83 (0.01)
Style Informal_to_formal 0.57 (0.01) 0.48 (0.02) 0.54 (0.09) 0.44 (0.05) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00) 0.07 (0.07) 0.29 (0.07) 0.25 (0.00)
median score 0.49 0.66 0.64 0.76 0.71
# best-performing tasks 7 8 8 15 14
Table 5: Average test performance (and standard deviations) across 3 random seeds comparing ACING versus APE, InstructZero, and INSTINCT. The bottom rows report the median score and total number of best-performing tasks for each method.
Appendix D ACING with Different Intrinsic (Action) Dimensions
In the main paper, we present results using actions with a dimension of $d^{\prime}=10$, following the setup of prior work. To evaluate the performance of ACING across different dimensionalities, we conducted experiments with $d^{\prime}\in\{5,10,20,40\}$, keeping other parameters fixed, for a budget of $165$. We report the test results over different tasks and dimensionalities for a fixed seed. The results, shown in Table 6, indicate that while the smallest dimension, $d^{\prime}=5$, recovered the best scores for some tasks, it generally has the lowest performance across most tasks. Furthermore, both $d^{\prime}=10$ and $d^{\prime}=20$ yield similar performance in terms of the number of best-performing tasks (9-10 tasks), indicating low sensitivity to this parameter. For the much larger dimension, $d^{\prime}=40$, the method achieved the highest number of best-performing tasks (15 tasks), demonstrating improved performance with increased dimensionality. Further increasing the dimensionality to $d^{\prime}=100$ can still yield high results, outperforming $d^{\prime}\in{5,10,20}$. However, while it remarkably outperformed $d^{\prime}=40$ in some tasks, such as the second word letter task, synonyms, and antonyms, it only achieved 14 best-performing tasks overall, indicating similar but slightly lower performance than $d^{\prime}=40$.
Category Task $d^{{}^{\prime}}=5$ $d^{{}^{\prime}}=10$ $d^{{}^{\prime}}=20$ $d^{{}^{\prime}}=40$ $d^{{}^{\prime}}=100$
(main paper)
Spelling Letters_list 1.00 1.00 0.98 1.00 1.00
First_word_letter 1.00 1.00 0.97 1.00 1.00
Second_word_letter 0.23 0.91 0.30 0.29 0.92
Morpho-Syntax Singular_to_plural 0.99 0.99 1.00 1.00 1.00
Active_to_passive 1.00 1.00 1.00 1.00 1.00
Negation 0.82 0.80 0.84 0.81 0.70
Lexical Semantics Antonyms 0.73 0.76 0.76 0.82 0.84
Synonyms 0.12 0.13 0.14 0.14 0.34
Word_unscrambling 0.53 0.54 0.49 0.55 0.43
Phonetics Rhymes 0.95 0.36 0.94 1.00 1.00
Numerical Sum 0.99 0.79 1.00 0.99 1.00
Diff 0.89 1.00 1.00 1.00 1.00
Knowledge Larger_animal 0.79 0.94 0.93 0.65 0.68
Periodic_elements 1.00 0.98 0.94 0.98 0.98
Cognitive Tasks Cause_and_effect 0.64 0.52 0.92 0.56 0.56
Common_concept 0.12 0.23 0.11 0.12 0.02
Object_counting 0.51 0.39 0.48 0.59 0.44
Odd_one_out 0.60 0.64 0.64 0.68 0.26
Orthography_starts_with 0.11 0.65 0.59 0.61 0.71
Taxonomy_animal 0.79 0.68 0.59 0.85 0.97
Auto_categorization 0.30 0.28 0.13 0.33 0.32
Word_sorting 0.55 0.69 0.74 0.69 0.48
CLUE Sentence_similarity 0.00 0.21 0.00 0.14 0.07
Sentiment 0.91 0.88 0.86 0.91 0.80
Translation Num_to_verbal 0.99 1.00 1.00 1.00 1.00
Translation_en-de 0.83 0.81 0.81 0.80 0.81
Translation_en-es 0.89 0.90 0.91 0.91 0.86
Translation_en-fr 0.84 0.84 0.86 0.88 0.73
Style Informal_to_formal 0.54 0.40 0.51 0.49 0.50
Coding Auto_debugging 0.25 0.25 0.25 0.25 0.25
# best-performing tasks 8 9 10 15 14
Table 6: Average ACING test performance for a fixed random seed (0) with different soft prompt dimensions $d^{{}^{\prime}}$. The bottom row report the total number of best-performing tasks.
Appendix E ACING with Different Number of Exemplars
In this section, we test ACING with a single exemplar, in contrast to the main results in the paper, which use five exemplars for ACING and all other benchmarks. For these experiments, we fix all hyperparameters as in the main paper and run tests with a budget of $165$. Intuitively, providing more exemplars to the language model should facilitate prompt learning, so five exemplars are expected to yield better prompts than a single exemplar. Our experiments, summarized in Table 7, support this intuition. The results show that using five exemplars leads to higher test scores, as reflected in a greater number of best-performing tasks and an increase in median test scores across tasks. However, it is notable that performance did not decrease drastically with only one exemplar, suggesting that a single exemplar is sufficient to achieve decent results. In fact, across several tasks and categories (e.g., phonetics, summation, morpho-syntax, and translation), a single exemplar achieves the same performance of using five exemplars, and even outperforms the use of five exemplars in certain tasks. Nevertheless, using a single exemplar resulted in lower performance mainly in more cognitively challenging tasks, which is understandable, as more complex tasks are likely to benefit from additional exemplars.
Category Task ACING ($|\mathcal{E}|=1$) ACING ($|\mathcal{E}|=5$)
(main paper)
Spelling Letters_list 1.00 (0.00) 1.00 (0.00)
First_word_letter 0.99 (0.01) 1.00 (0.00)
Second_word_letter 0.19 (0.09) 0.70 (0.15)
Morpho-Syntax Singular_to_plural 1.00 (0.00) 0.95 (0.03)
Active_to_passive 1.00 (0.00) 1.00 (0.00)
Negation 0.76 (0.08) 0.71 (0.06)
Lexical Semantics Antonyms 0.78 (0.05) 0.74 (0.01)
Synonyms 0.09 (0.03) 0.13 (0.02)
Word_unscrambling 0.41 (0.09) 0.50 (0.07)
Phonetics Rhymes 0.89 (0.08) 0.57 (0.31)
Numerical Sum 0.99 (0.01) 0.69 (0.31)
Diff 0.99 (0.01) 1.00 (0.00)
Knowledge Larger_animal 0.63 (0.17) 0.84 (0.07)
Periodic_elements 0.91 (0.08) 0.98 (0.00)
Cognitive Tasks Cause_and_effect 0.51 (0.08) 0.69 (0.15)
Common_concept 0.16 (0.11) 0.19 (0.05)
Object_counting 0.26 (0.06) 0.41 (0.03)
Odd_one_out 0.64 (0.02) 0.64 (0.00)
Orthography_starts_with 0.06 (0.05) 0.60 (0.12)
Taxonomy_animal 0.63 (0.06) 0.71 (0.02)
Auto_categorization 0.01 (0.01) 0.29 (0.04)
Word_sorting 0.70 (0.02) 0.70 (0.03)
CLUE Sentence_similarity 0.07 (0.05) 0.13 (0.07)
Sentiment 0.70 (0.12) 0.89 (0.01)
Translation Num_to_verbal 1.00 (0.00) 0.99 (0.01)
Translation_en-de 0.72 (0.11) 0.82 (0.01)
Translation_en-es 0.88 (0.00) 0.87 (0.02)
Translation_en-fr 0.16 (0.04) 0.83 (0.01)
Style Informal_to_formal 0.42 (0.05) 0.44 (0.05)
Coding Auto_debugging 0.25 (0.00) 0.25 (0.00)
median score 0.67 0.71
# best-performing tasks 12 23
Table 7: Average ACING test performance (and standard deviations) across 3 random seeds comparing 1 exemplar versus 5 exemplars. The bottom rows report the median score and total number of best-performing tasks.
Appendix F ACING with Different White-box models
[…]
Category Task ACING (Vicuna) ACING (WizardLM)
(main paper)
Spelling Letters_list 1.00 (0.00) 1.00 (0.00)
First_word_letter 1.00 (0.00) 1.00 (0.00)
Second_word_letter 0.70 (0.15) 0.36 (0.18)
Morpho-Syntax Singular_to_plural 0.95 (0.03) 0.99 (0.00)
Active_to_passive 1.00 (0.00) 1.00 (0.00)
Negation 0.71 (0.06) 0.83 (0.00)
Lexical Semantics Antonyms 0.74 (0.01) 0.81 (0.02)
Synonyms 0.13 (0.02) 0.12 (0.03)
Word_unscrambling 0.50 (0.07) 0.57 (0.05)
Phonetics Rhymes 0.57 (0.31) 0.97 (0.04)
Numerical Sum 0.69 (0.31) 1.00 (0.00)
Diff 1.00 (0.00) 1.00 (0.00)
Knowledge Larger_animal 0.84 (0.07) 0.94 (0.01)
Periodic_elements 0.98 (0.00) 0.97 (0.02)
Cognitive Tasks Cause_and_effect 0.69 (0.15) 0.76 (0.20)
Common_concept 0.19 (0.05) 0.21 (0.05)
Object_counting 0.41 (0.03) 0.46 (0.07)
Odd_one_out 0.64 (0.00) 0.56 (0.11)
Orthography_starts_with 0.60 (0.12) 0.62 (0.03)
Taxonomy_animal 0.71 (0.02) 0.60 (0.32)
Auto_categorization 0.29 (0.04) 0.35 (0.03)
Word_sorting 0.70 (0.03) 0.61 (0.02)
CLUE Sentence_similarity 0.13 (0.07) 0.22 (0.04)
Sentiment 0.89 (0.01) 0.90 (0.02)
Translation Num_to_verbal 0.99 (0.01) 1.00 (0.00)
Translation_en-de 0.82 (0.01) 0.81 (0.01)
Translation_en-es 0.87 (0.02) 0.61 (0.38)
Translation_en-fr 0.83 (0.01) 0.83 (0.05)
Style Informal_to_formal 0.44 (0.05) 0.32 (0.19)
Coding Auto_debugging 0.25 (0.00) 0.38 (0.10)
median score 0.71 0.79
# best-performing tasks 14 21
Table 8: Average ACING test performance (and standard deviations) across 3 random seeds using Vicuna and WizardLM as white-box models. The bottom rows report the median score and total number of best-performing tasks.
Appendix G Demonstrations with Human Instructions
Category Task Demonstration Human Instruction Honovich et al. (2023) Human Score Our Score
Spelling First_word_letter cat $\rightarrow$ c Extract the first letter of the input word. 1.00 1.00
Second_word_letter cat $\rightarrow$ a Extract the second letter of the input word. 0.96 0.59
Letters_list cat $\rightarrow$ c a t Break the input word into letters, separated by spaces. 1.00 1.00
Morpho- syntax Singular_to_plural cat $\rightarrow$ cats Convert the input word to its plural form. 1.00 1.00
Active_to_passive The artist introduced the scientist. $\rightarrow$ The scientist was introduced by the artist. Write the input sentence in passive form. 1.00 1.00
Syntax Negation Time is finite $\rightarrow$ Time is not finite. Negate the input sentence. 0.81 0.82
Lexical Semantics Antonyms won $\rightarrow$ lost Write a word that means the opposite of the input word. 0.70 0.82
Synonyms alleged $\rightarrow$ supposed Write a word with a similar meaning to the input word. 0.14 0.13
Phonetics Rhymes sing $\rightarrow$ ring Write a word that rhymes with the input word. 0.61 1.00
Knowledge Larger_animal koala, snail $\rightarrow$ koala Write the larger of the two given animals. 0.94 0.94
Semantics Cause_and_effect Sentence 1: The soda went flat. Sentence 2: The bottle was left open. $\rightarrow$ The bottle was left open. Find which of the two given cause and effect sentences is the cause. 0.97 0.92
Common_concept guitars, pendulums, neutrinos $\rightarrow$ involve oscillations. Find a common characteristic for the given objects. 0.11 0.11
Style Informal_to_formal Please call once you get there $\rightarrow$ Please call upon your arrival. Rephrase the sentence in formal language. 0.63 0.5
Numerical Sum 22 10 $\rightarrow$ 32 Sum the two given numbers. 1.00 1.00
Diff 32 22 $\rightarrow$ 10 Subtract the second number from the first. 1.00 1.00
Num_to_Verbal 26 $\rightarrow$ twenty-six Write the number in English words. 1.00 1.00
Multi- lingual Translation_en-de game $\rightarrow$ Spiel Translate the word into German. 0.81 0.84
Translation_en-es game $\rightarrow$ juego Translate the word into Spanish. 0.89 0.88
Translation_en-fr game $\rightarrow$ jeu Translate the word into French. 0.86 0.87
GLUE Sentiment The film is small in scope, yet perfectly formed. $\rightarrow$ positive Determine whether a movie review is positive or negative. 0.89 0.90
Sentence_similarity Sentence 1: A man is smoking. Sentence 2: A man is skating. $\rightarrow$ 0 - definitely not Rate the semantic similarity of two input sentences on a scale of 0 - definitely not to 5 - perfectly. 0.00 0.21
median score 0.89 0.90
average score 0.78 0.79
# best-performing tasks 14 16
Table 9: Classified tasks into categories from the instruction-induction datasets. For each task, we provide a corresponding demonstration, with $\rightarrow$ separating the input from the output, along with its respective human instruction as proposed in Honovich et al. (2023). We tested these instructions, report their test scores, and compare them to our best test scores using ACING with Vicuna-13B as the white-box LLM.
Appendix H Our Best Learned Instructions
Task Best instruction Test Score
active_to_passive Change the input to match the output, but the output is already in the passive voice 1.00
antonyms Take a word and change it to its opposite 0.82
auto_categorization Match the input to the output, and the answer is:$Input$: Nature Nanotechnology, Annual Review of Biochemistry, and The Lancet Neurology $Output$: top journals $Input$: Jeans, Tops, and Suits $Output$: Apparel 0.34
auto_debugging Input the code into a Python interpreter and observe the output 0.375
cause_and_effect Find the sentence that is the cause of the effect in the pair of sentences 0.92
common_concept Make a connection between the input and output, but the connection is not clear 0.11
diff Find the difference between the two numbers 1.00
first_word_letter Create a function that takes a string as input and returns the first letter of the first word in the string 1.00
informal_to_formal Convert the input into output using the same word order and with the same meaning 0.50
larger_animal Create a program that takes two animals as input and outputs the animal that is bigger 0.94
letters_list Input the word "year" and the output was "y e a r" 1.00
negation Flip the truth value of the statements in the input 0.82
num_to_verbal Convert numbers to words 1.00
object_counting Provide a number that represents how many items are in the input 0.55
odd_one_out Find the word that does not belong in each group based on the given words 0.64
orthography_starts_with Find a word in the text that starts with the letter provided and to output that word 0.71
periodic_elements Find the name of the element based on its atomic number 1.00
rhymes Input the word that the program thought I was inputting and then output the word that program thought I was inputting 1.00
second_word_letter Output the first letter of each word that is a vowel (a, e, i, o, u) Example: Input: year Output: e Input: trust Output: r Input: qualification Output: u Input: defendant Output: 0.59
sentence_similarity Find a sentence pair that is probably not similar, and the output is 3 - probably 0.21
sentiment Classify each input as positive or negative based on the assessment of the corresponding movie 0.90
singular_to_plural Add the suffix -s to the end of the word to make it plural 1.00
sum Find the sum of the two numbers 1.00
synonyms Input a word that is a synonym for the word that was output 0.13
taxonomy_animal Make the AI generate a sequence of animals based on the input provided 0.75
translation_en-de Provide a translation for each word in the English text into German 0.84
translation_en-es Translate the words from English to Spanish, but I noticed that some of the translations are not accurate 0.88
translation_en-fr Create a program that would take an English word as input and output its French equivalent 0.87
word_sorting Output the words in alphabetical order, but the output is not in alphabetical order 0.73
word_unscrambling Convert the input to a word that is a common English word 0.62
Table 10: The best instruction discovered by ACING for all the 30 instruction-induction tasks using with Vicuna-13B as the white-box LLM.